### v2

@@ 4.3 Specific cases
Item 2: completely analogous to Item 3, observing that $\tilde{\Gamma}_{F}=\operatorname{diag}(\tilde{\gamma}_{f}M_{z},0)$.
Theorem 4.17 (PDPS with adjoint mismatch).
Assume the setup of Example 4.2 with $\tau\sigma\|K\|^{2}\leq 1$ and, for simplicity, $f=0$ and Hilbert $Z$ and $Y$. Suppose $\operatorname{dom}h_{*}$ is bounded, and that $g$ and $h_{*}$ are, respectively, $\gamma_{g}$- and $\gamma_{h_{*}}$-strongly convex for some $\gamma_{g}>0$ and $\gamma_{h_{*}}\geq 0$. Let $\gamma:=\min\{\gamma_{g}\tau/4,\gamma_{h_{*}}\sigma/2\}$. In the PDPS (30), not able to compute $K^{*}$, replace it with a “mismatched” adjoint $K^{*\approx}$. Then, for any $\bar{x}\in Z\times Y$ and $p\in(1,1+2\gamma]$, Section 4.2Item (a) holds with $\breve{\Lambda}=0$, $\Omega_{\bar{x}}=Z\times Y$, $\delta=\infty$, $r_{p}\leq\varepsilon/(1-p)$, and
Proof 4.18.
With $M$, $G$, and $F$ given by Example 4.2, the abstract algorithm (28) reads
$-M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}=x_{k+1}^{*}+((K^{*\approx}-K^{*})y^{k},0)\quad\text{for a}\quad x_{k+1}^{*}\in H(x^{k+1}).$
Here $H$ is defined in (27). Using Lemma 4.3 in the final step, we estimate
~$\langle\tilde{\partial}_{k+1}-H(\hat{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$ =~$=\langle\tilde{\partial}_{k+1}-x_{k+1}^{*}|x^{k+1}-\bar{x}\rangle_{X^{*},X}+\langle x_{k+1}^{*}-H(\hat{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$
$\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\varepsilon_{k}(\bar{x}).$
Therefore, (33) holds with the stated choices. Moreover, we have $\sum_{k=0}^{N-1}p^{k-N}\leq 1/(p-1)$ for any $p\in(1,1+2\gamma]$, verifying (35) and consequently Section 4.2Item (a).
Remark 4.19 (Stochastic optimisation methods).
These can be approached through lifting: we take $X$ as a space of random variables on a space $\mathscr{X}$, set $F(x)=\mathbb{E}[\mathscr{F}\circ x]$ for a function $\mathscr{F}$ on $\mathscr{X}$, $Mx:=\mathbb{E}^{*}\mathscr{M}\mathbb{E}[x]$ for $\mathscr{M}\in\mathbb{L}(\mathscr{X};\mathscr{X}^{*})$, etc., where $\mathbb{E}$ is the expectation. Without inexactness, this produces parallel copies of the very same optimisation method for every random event $\omega$. We then model the random choices made on every step through inexactness, subject to Sections 4.2, 4.2 and 4.2.
We first show the potentially global convergence of subdifferentials; see Remark 4.8. When $\Xi=0$, this could be followed by the Kurdyka–Łojasiewicz property to show function value convergence, and, afterwards, either by a growth condition or, in finite dimensions, a finite-length argument based on (43) and [1, proof of Lemma 2.6] to show iterate convergence. As the property can easily be verified only in finite dimensions (for semi-algebraic functions), we prefer a more direct approach.
Theorem 4.20.
If Section 4.2 holds, then $x^{k}\in\Omega$ and
If, moreover, Section 4.2 holds, then also $\inf_{x^{*}\in H(x^{k+1})}\|x^{*}\|_{X^{*}}\to 0$
Proof 4.21.
By the implicit algorithm (28), the properties of Fenchel conjugates (e.g., [7, Lemma 5.7]) and $-M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}\in\partial\left(\frac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\right)(x^{k+1}-x^{k})$, we have
If $\{x^{j}\}_{j=0}^{N-1}\subset\Omega$, Section 4.2Item (i) thus yields for all $k=0,\ldots,N-1$ that
Summing over all such $k$, and using Section 4.2Item (ii), it follows
From Section 4.2Item (iii), it now follows that $x^{N}\in\Omega$. Since, by the same assumption, $x^{0}\in\Omega$, induction establishes (43) and $x^{k}\in\Omega$ for all $k\in\mathbb{N}$. Using Section 4.2Item (iv) in (46), we, moreover, deduce $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ and $(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1})\to 0$. Let $c\geq\|M\|_{\mathbb{L}(X;X^{*})}$. By $\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\leq c\|\,\boldsymbol{\cdot}\,\|_{X}^{2}$ and the properties of conjugates (e.g., [7, Lemmas 5.4 and 5.7]),
Thus also ~$\|\tilde{\partial}_{k+1}\|_{X^{*}}\to 0$. Section 4.2 proves that $\inf_{x^{*}\in H(x^{k+1})}\|\tilde{\partial}_{k+1}-x^{*}\|_{X^{*}}\to 0$. Hence an application of the triangle inequality establishes $\inf_{x^{*}\in H(x^{k+1})}\|x^{*}\|_{X^{*}}\to 0$.
Example 4.22 (Forward-backward splitting).
For the (inexact) forward-backward splitting of Examples 4.1, 4.13 and 4.14, the theorem establishes the monotonicity of function values, as well as the convergence of subdifferentials to zero, $\inf_{x^{*}\in\partial G(x^{k+1})}\|F^{\prime}(x^{k+1})+x^{*}\|\to 0$.
Remark 4.23 (Bregman divergences).
The argument of Theorem 4.20 extends to algorithms where the $M$-seminorm for $M\in\mathbb{L}(X;X^{*})$ is replaced by a Bregman divergence $B_{M}$ generated by some convex $M:X\to\overline{\mathbb{R}}$. Recalling the definition
in the algorithm (28), we would impose ~$\tilde{\partial}_{k+1}=\omega^{k+1}-\omega^{k}\in\partial_{2}B_{M}^{\omega^{k}}(x^{k},x^{k+1})$ for a given $\omega^{k}\in\partial M(x^{k})$ and some $\omega^{k+1}\in\partial M(x^{k+1})$. Using the Fenchel–Young identity, we could then replace (44) by
Thus repeating the arguments of the theorem would establish both $B_{M}^{\omega^{k}}(x^{k},x^{k+1})\to 0$ as well as $B_{M^{*}}^{x^{k+1}}(\omega^{k+1},\omega^{k})=B_{M^{*}}^{x^{k+1}}(\omega^{k}+\tilde{\partial}_{k+1},\omega^{k})\to 0$. A variant of Section 4.2 could then establish a form of convergence for $H(x^{k+1})$.
4.5 Non-escape, quasi-Féjer monotonicity, linear convergence
The next lemma is essential for all our strong convergence results. The proof is standard; see, e.g., [7, Chapter 15] for the case $\varepsilon_{k}(\bar{x})=0$ and $\Xi=0$. Observe that (48) with the triangle inequality may be used to again prove Section 2.1Item (i) for multilevel methods.
Lemma 4.24.
Suppose Section 4.2 holds at $\bar{x}\in X$. Then $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for all $k\in\mathbb{N}$, and the sequence is ($p$-strongly) quasi-Féjer, i.e.,
Proof 4.25.
We first treat Section 4.2 option Item (a). Fix $N\in\mathbb{N}$ and suppose $\{x^{j}\}_{j=0}^{N-1}\subset\Omega_{\bar{x}}$. Observe that $\langle\Xi x|x\rangle=0$ for all $x\in X$ by the skew-adjointness of $\Xi$. Since $0\in H(\bar{x})$, using (33) in the implicit algorithm (28), we thus get
for all $k\in\{0,\ldots,N-1\}$. By $\breve{\Lambda}\leq(1-\eta)M$ and the Pythagoras’ identity (see [34, (2.3)])
$\langle M(x-z)|x-\bar{x}\rangle=\frac{1}{2}\|x-z\|_{M}^{2}+\frac{1}{2}\|x-\bar{x}\|_{M}^{2}-\frac{1}{2}\|z-\bar{x}\|_{M}^{2}\quad(x,z,\bar{x}\in X),$
we obtain
Hence $x^{N}\in\mathbb{O}_{M}(\bar{x},\delta)$. Since $x^{0}\in\Omega_{\bar{x}}$ by Section 4.2, an inductive argument shows that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for all $k\in\mathbb{N}$, justifying the above steps. Finally, (49) shows (48), while $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ follows from (51) and $\eta>0$.
Regarding option Section 4.2Item (b), arguing as above with (34) in place of (33), we get in place of (49) the estimate
Using $\inf_{x\in\mathbb{O}_{M}(\delta,\bar{x})}\mathcal{G}(x;\bar{x})\geq 0$, we proceed (with $1+\gamma\geq p$) as in option Item (a) to establish (51), and from there onwards.
@@ 4.5 Non-escape, quasi-Féjer monotonicity, linear convergence
Suppose Section 4.2 holds at $\bar{x}\in X$ with $p>1$ and the inequality in (35) strengthened to
Then $\|x^{N}-\bar{x}\|_{M}^{2}\to 0$ at the rate $O(p^{-N})$.
Example 4.27.
From the proof of Theorem 4.11, it is clear that (53) holds for the exact and inexact forward backwards methods of Examples 4.13 and 4.14. Since Theorem 4.15 is based on Theorem 4.11, it also holds for exact and inexact primal-dual methods. For $p>1$, sufficient second-order growth (strong convexity), is still required.
4.6 Local convergence of function values
We now proceed to function values and duality gaps. The idea of possibly assuming both Section 4.2Item (a) and a relaxed version of Item (b), as an alternative to just the latter, is to be able to study descent at non-minimising critical points. For simplicity, we only treat sublinear convergence.
Theorem 4.28.
Suppose Section 4.2 holds at $\bar{x}\in X$ and, for a non-empty set $\hat{X}\subset X$, Eq. 34 holds for all $\hat{x}\in\hat{X}$ with $\breve{\Lambda}=\breve{\Lambda}_{\hat{x}}\leq M$, $\gamma=\gamma_{\hat{x}}\geq 0$, and $\Omega_{\hat{x}}\supset\mathbb{O}_{M}(\bar{x},\delta)$. Then
If $\Xi=0$ and Section 4.2 holds [^3] Since the proof of the present Theorem 4.28 shows that $x^{k}\in\mathbb{O}_{M}(\hat{x},\delta)$ for all $k\in\mathbb{N}$, to prove the required (43), it would be enough to assume that just Section 4.2Item (i) holds with $\Omega\supset\mathbb{O}_{M}(\hat{x},\delta)$., then, for all $N\in\mathbb{N}$,
(55) $[F+G](x^{N})\leq\inf_{\hat{x}\in\hat{X}}[F+G](\hat{x})+\sup_{\hat{x}\in\hat{X}}\left(\frac{1}{2N}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\left(\frac{1}{N}\varepsilon_{\mathrm{desc},k}(\hat{x})+\frac{k+1}{N}\varepsilon_{\mathrm{desc},k}\right)\right).$
Proof 4.29.
Lemma 4.24 shows for all $k\in\mathbb{N}$ that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\bigcap_{\hat{x}\in\hat{X}}\Omega_{\hat{x}}$. Hence, for any $\hat{x}\in\hat{X}$, we may follow the proof of the lemma for case Item (b) of Section 4.2 to establish (52) for $\bar{x}=\hat{x}$. To reach this point, the assumption $\inf_{x\in\mathbb{O}_{M}(\delta,\bar{x})}\mathcal{G}(x;\bar{x})\geq 0$ was not yet needed. Now, summing (52) over $k=0,\ldots,N-1$, we obtain
Taking the supremum over $\hat{x}\in\hat{X}$, this establishes (54).
Suppose then that $\Xi=0$ and Section 4.2 holds. Theorem 4.20 now establishes (43), i.e., the quasi-monotonicity $[F+G](x^{k+1})\leq[F+G](x^{k})+\varepsilon_{\mathrm{desc},k}.$ Repeatedly using this and $\mathcal{G}(x^{k+1};\hat{x})=[F+G](x^{k+1})-[F+G](\hat{x})$ in (56), and dividing by $N$, we obtain (55).
We next specialise the result to the PDPS of Example 4.2. Besides inexactness, as a novelty compared to [5, 6, 18, 11], subject to $h_{*}$ having a bounded domain, we get an estimate on the convex envelope of the objective, i.e., the Fenchel biconjugate. In non-reflexive spaces, we define the latter as a function in $X$ instead of $X^{**}$ by taking first the conjugate and then the equivalently defined preconjugate: $h^{**}:=(h_{*})_{*}$.
Corollary 4.30.
@@ 4.6 Local convergence of function values
Assume the setup of Examples 4.2 and 4.1 for some $\tau,\sigma,\lambda>0$, as well as that Theorem 4.15Items 2 and 3 hold for $p=1$ at some $\bar{z}\in Z$ with $\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\subset\Omega_{\bar{z}}$ for some $\delta_{z}>0$. Also suppose that that $\operatorname{dom}h_{*}$ is bounded, $0\in H(\bar{x})$ for some $\bar{x}\in\{\bar{z}\}\times\operatorname{dom}h_{*}$, and that the condition (39) on the initial iterate holds. Then, for the ergodic iterates $\tilde{z}^{N}:=\frac{1}{N}\sum_{k=0}^{N-1}z^{k}$, for all $N\in\mathbb{N}$, we have
$[f+g+h\circ K]^{**}(\tilde{z}^{N})\leq[f+g+h\circ K](\bar{z})+\sup_{\hat{y}\in\operatorname{dom}h_{*}}\frac{1}{2N}\|(z^{0},y^{0})-(\bar{z},\hat{y})\|_{M}^{2}+\frac{\sum_{k=0}^{N-1}e_{1,k}(z^{k+1})}{2\tilde{\gamma}\lambda N}.$
Here $[f+g+h\circ K](\bar{z})=[f+g+h\circ K]^{**}(\bar{z})$ if $\bar{z}$ is a global minimiser of $f+g+h\circ K$.
Proof 4.31.
Theorem 4.15Item 2 proves Section 4.2 option Item (a) at $\bar{x}$. Likewise, Theorem 4.15Item 3 shows (34) and $\Omega_{\hat{x}}:=\Omega_{\bar{z}}\times\operatorname{dom}h_{*}\supset\mathbb{O}_{M}(\bar{x},\delta)$ at any $\hat{x}\in\hat{X}:=\{\bar{z}\}\times\operatorname{dom}h_{*}$. Theorem 4.28 now establishes (54), whose left-hand-side we still have to estimate.
@@
With the expression of Example 4.7 for the gap, we expand and estimate using the definition of the Fenchel (bi)conjugate and $h^{**}=h$ as well as $[f+g]^{**}\leq f+g$ that
$\mathcal{G}(x^{k+1};\bar{x})$ $=([f+g](z^{k+1})+\langle Kz^{k+1}|\hat{y}\rangle-h_{*}(\hat{y}))-([f+g](\bar{z})+\langle K\bar{z}|y^{k+1}\rangle-h_{*}(y^{k+1}))$
$\geq\bigl{(}[f+g]^{**}(z^{k+1})+\langle Kz^{k+1}|\hat{y}\rangle-h_{*}(\hat{y})\bigr{)}-N[f+g+h\circ K](\bar{z}).$
@@
$\sup_{\hat{y}\in\operatorname{dom}h_{*}}\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\bar{x})\geq N[(f+g)^{**}+h\circ K](\tilde{z}^{N})-N[f+g+h\circ K](\bar{z}).$

### v3

@@ 6.3 Primal-dual proximal splitting
Item˜2: completely analogous to Item˜3, observing that $\tilde{\Gamma}_{F}=\operatorname{diag}(\tilde{\gamma}_{f}M_{z},0)$.
Remark 6.15 (Choice of the parameters).
If $f^{\prime}$ is locally $\gamma_{f}$-strongly monotone, the conditions Eqs.˜58 and 59 can be ensured by taking the free parameters $\tilde{\gamma},\zeta,\beta>0$ sufficiently small (to ensure Eqs.˜58a and 59a), and then $\lambda>0$ sufficiently large (to ensure Eqs.˜58b and 59b). Through the basic step length conditions Section˜6.3, a large $\lambda$ will force the primal step length $\tau>0$ to small, and allow the dual step length $\sigma>0$ to be correspondingly large. The initial iterate should also satisfy the locality bound (57), where both $\lambda$ and $\tau$ appear, the latter through $M$ defined in (25b). For Section˜4.2 only the basic step length conditions are required besides the tracking estimates on $f$.
Now that we have provided step length and growth conditions that prove LABEL:, 4.2, 4.2 and 4.2 for single-loop PDPS for bilevel problems, we can use Theorems˜4.6, 4.13 and 4.16 to prove different forms of convergence. In fact, further specialising Theorem˜4.13 to the PDPS, besides inexactness, as a novelty compared to [5, 6, 20, 12], subject to $h_{*}$ having a bounded domain, we get an estimate on the convex envelope of the objective, i.e., the Fenchel biconjugate. In non-reflexive spaces, we define the latter as a function in $X$ instead of $X^{**}$ by taking first the conjugate and then the equivalently defined preconjugate: $h^{**}:=(h_{*})_{*}$.
@@
Assume the setup of Eqs.˜54 and 6.3 for some $\tau,\sigma,\lambda>0$, as well as that Theorem˜6.13Items˜2 and 3 hold for $p=1$ at some $\bar{z}\in Z$ with $\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\subset\Omega_{\bar{z}}$ for some $\delta_{z}>0$. Also suppose that that $\operatorname{dom}h_{*}$ is bounded, $0\in H(\bar{x})$ for some $\bar{x}\in\{\bar{z}\}\times\operatorname{dom}h_{*}$, and that the condition (57) on the initial iterate holds. Then, for the ergodic iterates $\tilde{z}^{N}:=\frac{1}{N}\sum_{k=0}^{N-1}z^{k}$, for all $N\in\mathbb{N}$, we have
$f+g+h\circ K]^{**}(\tilde{z}^{N})\leq[f+g+h\circ K](\bar{z})+\sup_{\hat{y}\in\operatorname{dom}h_{*}}\frac{1}{2N}\|(z^{0},y^{0})-(\bar{z},\hat{y})\|_{M}^{2}+\frac{\sum_{k=0}^{N-1}e_{1,k}}{2\tilde{\gamma}N}.$
Here $f+g+h\circ K](\bar{z})=[f+g+h\circ K]^{**}(\bar{z})$ if $\bar{z}$ is a global minimiser of $f+g+h\circ K$.
Proof 6.17.
[…]
(61) $\sup_{\hat{x}\in\hat{X}}\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\hat{x})\leq\sup_{\hat{x}\in\hat{X}}\left(\frac{1}{2}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\frac{e_{1,k}}{2\tilde{\gamma}}\right)\quad\text{for all}\quad N\in\mathbb{N}.$
@@
Let $\hat{x}=(\bar{z},\hat{y})\in\hat{X}$. With the expression of Example˜4.3 for the gap, we expand and estimate using the definition of the Fenchel (bi)conjugate and $h^{**}=h$ as well as $f+g]^{**}\leq f+g$ that
[…]
$\geq\bigl([f+g]^{**}(z^{k+1})+\langle Kz^{k+1}|\hat{y}\rangle-h_{*}(\hat{y})\bigr)-[f+g+h\circ K](\bar{z}).$
@@
$\sup_{\hat{y}\in\operatorname{dom}h_{*}}\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\hat{x})\geq N[(f+g)^{**}+h\circ K](\tilde{z}^{N})-N[f+g+h\circ K](\bar{z}).$